Using 26 letters (non-case-sensitive), how many password can be made if you can repeat letters for a password of 5 characters?

Question 10 options:

11,881,376

65,780

7,893,600

358,800
Answer #1

26 letters (non-case-sensitive) means only 26 letters.

Each password of 5 characters.

can repeat letters for a password mean one word can come again or repetitive in password.

with 26 letters 5 character password & with repetition

we can made = 26 ^ 5 passwords

= 11881376

Answer:- 11881376

(Note :- In no repeatation of characters

, 5! * (26C1 * 25C1 * 24C1 * 23C1*22C1) = 120*(26*25*24*23*22) =

947232000

passwords are possible.